TVM

You need a 30-year, fixed-rate mortgage to buy a new home for $220,000. Your mortgage bank will lend you the money at a 6.8 percent APR for this 360-month loan. However, you can afford monthly payments of only $1,100, so you offer to pay off any remaining loan balance at the end of the loan in the form of a single balloon payment. How large will this balloon payment have to be for you to keep your monthly payments at $1,100?
Answer #1

FV of $220,000 after 30 years:

FV= 220000*(1+6,8%)^30=1,682,219

FV of the $1,100 monthly payment at the end of 30 years

image.png

=> FV= 1,290,194

The balloon payment = 1,682,219-1,290,194= $392025


answered by: trang trần